Let X be a Bernoulli random variable (Hint: Special case of Binomial Distribution)

a. Compute E(X2)

b. Show that V(X) is p(1-p)

c. Compute E(X79)


1
Expert's answer
2021-09-14T00:58:08-0400

Solution:

"X \sim \operatorname{Bernoulli}(p), 0 \leq p \leq 1"

So, "P(X=x)= \begin{cases}1-p & \text { if } x=0 \\ p & \text { if } x=1\end{cases}"  

Now the expected value of X:

 "\begin{aligned}

E(X) &=0 \cdot P(X=0)+1 \cdot P(X=1) \\

&=0 \times(1-p)+1 \times p \\

&=p

\end{aligned}"  

(a):

 "\begin{aligned}

E\left(X^{2}\right) &=0^{2} \cdot P(X=0)+1^{2} \cdot P(X=1) \\

&=0 \times(1-p)+1 \times p \\

&=p

\end{aligned}"  

(b):

 "\begin{aligned}

\operatorname{Var}(X) &=E\left(X^{2}\right)-E^{2}(X) \\

&=p-p^{2} \\

&=p(1-p)

\end{aligned}"

(c):

"\begin{aligned}

E\left(X^{79}\right) &=0^{79} \cdot P(X=0)+1^{79} \cdot P(X=1) \\

&=0 \times(1-p)+1 \times p \\

&=p

\end{aligned}"